ganeshie8 (ganeshie8):

show that \[\large p | (x^2+1) \text{ has a solution } \iff p = 4k+1\] \(p\) is an odd prime

OpenStudy (anonymous):

Well, let's see. we know that p is a prime number, that means it is odd (not even). But as you say it is defined as \(x^2 + 1\). That means \(x\) has to be an even integer, because \(x^2\) has to be an even integer. Only that way adding 1 to it would end up with an odd integer which is a requirement for a prime number. That means \(\frac{1}{2}x\) is an integer as well. So now we can say \(k = \bigg( \frac{1}{2}x\bigg)^2 = \frac{1}{4}x^2\) So now for any such nubmer we have a 'k' which satisfies $$ 4k + 1 = 4 \cdot \bigg( \frac{1}{4}x^2 \bigg) + 1 = x^2 + 1= p$$

OpenStudy (anonymous):

Well thanks. I didn't show the other way around, but it's very similar =)

OpenStudy (thomas5267):

Wow. I don't even understand the question...

OpenStudy (anonymous):

hmm is it Lagrange theorem ?

ganeshie8 (ganeshie8):

Yes... that works i guess

OpenStudy (anonymous):

\(x^2+1\equiv 0 \mod p\\ gcd(1,p)=1 ~~and~ gcd(4,p)=1\\ ~thus ~ \\x^2+1\equiv 4(x^2+1) \equiv 0 \mod p \\ 4x^2\equiv -4 \mod p \\let~ y=2x \\ (2x)^2\equiv -4 \mod p \\ y^2\equiv -4 \mod p \\ \text{let y_0 be a solution } \\ y_0\equiv -4 \mod p\\ y_0+4=kp\\ y_0=kp-4 \\ 2x\equiv y_0-p \mod p\) hmm dont feel this ends it good, i'll make lunch brb

OpenStudy (anonymous):

and i have something with @pitamar solution x need not to be odd hmm see 3|6 and 6 is even

ganeshie8 (ganeshie8):

I see you're using quadratic residues... that works perfectly, we need to tweak it a bit more ... Here is another proof with a big flaw : By Fermat's little theorem we have \(\large x^{p-1}\equiv 1\pmod{p}\tag{1}\) \(\large x^2\equiv -1\pmod{p}\) squaring both sides \(\large x^4\equiv 1\pmod{p}\tag{2}\) from (1) and (2) \[4|(p-1) \implies p = 4k+1\]

OpenStudy (anonymous):

eh that was so quick i dint even got the chance to think xD

OpenStudy (anonymous):

i was trying to use lagrange thingy

ganeshie8 (ganeshie8):

lagrange theorem gives you solution in one line

OpenStudy (anonymous):

haha yeah it only says blah blah have solution iff p=4k+1 =) so thought of thinking its proof

ganeshie8 (ganeshie8):

discrete logarithm is much better : \[x^2\equiv -1 \pmod{p}\] take discrete logarithm both sides : \[2 \log x \equiv \log (-1) \pmod{p-1}\] \[2 \log x \equiv \frac{p-1}{2} \pmod{p-1}\] for this congruence to admit a solution, we need \(\gcd(2, p-1) | \frac{p-1}{2}\)

ganeshie8 (ganeshie8):

but gcd(2, p-1) is 2 itself so we have \(2 | \dfrac{p-1}{2} \implies p = 4k+1\)

ganeshie8 (ganeshie8):

wanna see the proof using lagragne thm ?

OpenStudy (anonymous):

sure i forgot these stuff =)

OpenStudy (kainui):

I keep seeing these but I don't ever seem to understand them even after reading and sort of seeing an explanation for some reason, what's wrong with me lol

ganeshie8 (ganeshie8):

Another proof w/o appealing directly to lagrange thm : \(\large (\Leftarrow )\) suppose \(p = 4k+1\) and consider below congruence \(x^4\equiv 1 \pmod{p}\) factoring gives \((x^2+1)(x^2-1)\equiv 0 \pmod{p}\) because \(p\) is a prime it follows that \(x^2+1 \equiv 0 \pmod p\) or \(x^2-1\equiv 0 \pmod{p}\) letting \(x\) be primitive root, we see that \(x^2-1\not \equiv 0 \pmod{p}\) that yields \(x^2+1 \equiv 0 \pmod p\) \(\blacksquare\)

ganeshie8 (ganeshie8):

Maybe I'll let you work the proof using lagrange thm :)

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

Lagrange `if p is prime and f(x) polynomial with leading coefficient a such that` `gcd(a,p)=1 (great power is n)` `and` \( f(x)\equiv 0 \mod p\) `has at most n incongruent solution modulo p` so \(x^2+1\equiv p\) has 2 solution hehehe so far i tried previous stuff and got stuck

ganeshie8 (ganeshie8):

Oh that can be pretty lengthy.. we can use below version of lagrange thm : \(x^k\equiv a \pmod p\) has a solution \(\iff\) \(a^{(p-1)/d}\equiv 1\pmod{p}\) where \(d = \gcd(k,p-1)\)

OpenStudy (anonymous):

got it !!

OpenStudy (anonymous):

x^2=-1 mod p -1^(p-1)/2=1 mod p p-1 = 4k ( we need -1^b such that b is even) p=4k+1 x'D

OpenStudy (anonymous):

like p-1/2= 2k to make -1^something=1 p-1=4k

ganeshie8 (ganeshie8):

thats it! p = 4k+3 is not possible because then (p-1)/2 becomes odd

OpenStudy (anonymous):

thats really nice question

ganeshie8 (ganeshie8):

@Kainui I think all these theorems are based around discrete logarithm

ganeshie8 (ganeshie8):

we might be able to prove all the stated thms in this thread using just the definition of discrete logarithm http://en.wikipedia.org/wiki/Discrete_logarithm

OpenStudy (anonymous):

haha abstract algebra =)

OpenStudy (anonymous):

i thought that at first i couldnt memorize what it was exactly special with the pre question

OpenStudy (anonymous):

see a^k have all modulo classes of p :| and thats it QED!!